OpenStudy (anonymous):

I was wondering if I could solve a differential equation from my textbook in Chemical kinetics but I wasn't sure exactly how to tackle it - and if it's easy to solve on paper at all. Here is my equation:

OpenStudy (anonymous):

\[\frac{ dB }{ dt }=k1*A0*\exp(-k1*t) - k2*B\]

OpenStudy (anonymous):

I can't separate the variables, so I guess there is some other trick to get the result.

OpenStudy (anonymous):

This is a linear differential equation in B and t

OpenStudy (anonymous):

Do you know how to solve them?

OpenStudy (anonymous):

The only method we've used so far in our Chemistry classes is to separate the variables.

ganeshie8 (ganeshie8):

today u will be learning another trick to solve it then :)

ganeshie8 (ganeshie8):

familiar wid 'product rule' of differentiation right ?

OpenStudy (anonymous):

oh, thank you :)

OpenStudy (anonymous):

A linear differential equation is of the form y' + k(t)y = m(t) this corresponds to that form doesn't it?

OpenStudy (anonymous):

I am, yeah

OpenStudy (anonymous):

We will be using the product rule to solve this equation

OpenStudy (anonymous):

okay, listening :)

OpenStudy (anonymous):

So your equation is of the form I just wrote above isn't it? B' + k2*B = a0 * k1 * exp(-k1*t)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To solve this kind of equation we multiply it with something we call an integrating factor. After that we can use the product rule simplify the equation. In the form y' + k(t)y = m(y) the integrating factor is exp(integral(-k(t)dt)) okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

In your case the factor is exp(integral(k2dt)) = exp(k2*t) yes?

OpenStudy (anonymous):

ah, yeah, sorry I'm that slow

OpenStudy (anonymous):

Hey never mind.. Now please multiply the equation by that integrating factor and write it here

OpenStudy (anonymous):

just a question - so, in the equation we have k(t)y to correspond to k2B... and in the integrating factor we have the integral of k(t)... when k2 is just a constant?

OpenStudy (anonymous):

in this case -k2 is your k(t) and B is your y so IF is exp(k2*t)

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

ah, right - we assume that the k(t) has t to the power 0, so only a constant remains

OpenStudy (anonymous):

so, let me write the whole equation now

OpenStudy (anonymous):

we have \[\frac{ dB }{ dt }*e ^{-k2t} + k2Be ^{-k2t}=k1A0.e ^{-(k1+k2)t}\]

OpenStudy (anonymous):

The IF is exp(k2*t) correct that and then see if you can recognize the LHS as d(B*exp(k2*t))/dt

OpenStudy (anonymous):

oops, yeah, without the minus in the exponent

OpenStudy (anonymous):

\[\frac{ dB }{ dt }*e ^{k2t} + k2Be ^{k2t}=k1A0.e ^{(k2-k1)t}\]

OpenStudy (anonymous):

get the dt to the other side now you have B*exp(k2*t) = k1*a0*exp((k2-k1)t)dt \ isnt it?

OpenStudy (anonymous):

yes, but how about the t on the left side? or it's okay for it to stay there?

OpenStudy (anonymous):

can you solve it from here?

OpenStudy (anonymous):

let me write it on paper to see better

OpenStudy (anonymous):

Ugh, I thought I'd got it right but it turned out that there's some extra k2 in the answer I got compared to the one in the textbook. I'm attaching my solution. Help highly appreciated!

OpenStudy (turingtest):

when you integrated the lhs of the equation you pulled out a k_2 unnecessarily. Use the fundamental theorem of calculus; you are integrating a derivative here

OpenStudy (anonymous):

@TuringTest so in the left hand side, we assume e^k2*t is a constant and just take it out like this?

OpenStudy (turingtest):

you don't assume it's a constant, the integration is implicit in using the integrating factor

OpenStudy (turingtest):

lhs\[B'e^{k_2t}+K_2Be^{k_2t}=(Be^{k_2t})'\]so when you integrate it you just get\[Be^{k_2t}+C\]

OpenStudy (anonymous):

oh, got it now, thank you very much!

OpenStudy (turingtest):

welcome :)